### v1

# Atypical generic directions in Teichmüller space

Abstract.
Using work of Chaika–Masur–Wolf and Durham–Zalloum, we construct the first example of a sublinearly-Morse Teichmüller geodesic ray with minimal non-uniquely ergodic vertical foliation.
1. Introduction
There is delicate interplay between the geometry of the Teichmüller metric on the Teichmüller space, $\mathrm{Teich}(S)$, of a finite-type surface $S$ and its Thurston compactification, $\mathbb{P}\mathcal{MF}(S)$, by projectivized measured foliations. The crux of the complexity is that Teichmüller geodesics are defined by deformations of singular flat structures on $S$, while $\mathbb{P}\mathcal{MF}(S)$ is defined via hyperbolic geometry. This leads to an incongruity in how the internal geometry of $\mathrm{Teich}(S)$ is asymptotically encoded in $\mathbb{P}\mathcal{MF}(S)$. Notably, Lenzhen [Len08] proved that geodesic rays can have limit sets in $\mathbb{P}\mathcal{MF}(S)$ larger than a point (see also [CMW19, LLR18, BLMR20]).
In this note, we produce a concrete example highlighting how this tension manifests in the context of random walks of the mapping class group $\mathcal{MCG}(S)$ on $\mathrm{Teich}(S)$. For our purposes, the main question here is whether generic directions—namely geodesic rays tracked by sample paths of random walks—can be completely encoded via properties of the internal geometry of $\mathrm{Teich}(S)$.
In their breakthrough paper, Kaimanovich–Masur [KM96] proved that Thurston’s compactification of Teichmüller space is a topological model for the Poisson boundary for many random walks on the mapping class group, with sample paths tracking Teichmüller geodesic rays with uniquely ergodic vertical foliations. One upshot here is that the limit set in $\mathbb{P}\mathcal{MF}(S)$ of a generic direction in $\mathrm{Teich}(S)$ is a unique point, hence avoiding the above tension between $\mathrm{Teich}(S)$ and $\mathbb{P}\mathcal{MF}(S)$.
More recently, Gekhtman–Qing–Rafi [GQR22] proved that these tracking rays are sublinearly Morse [QR22, QRT24], a weak hyperbolicity condition which is controlled by the internal geometry of $\mathrm{Teich}(S)$ by work of Durham–Zalloum [DZ22]. In [GQR22], the authors proved that the set of accumulation points in $\mathbb{P}\mathcal{MF}(S)$ of all sublinearly Morse Teichmüller geodesic rays has full measure with respect to any stationary measure associated to a (sufficiently nice) random walk of $\mathcal{MCG}(S)$ on $\mathrm{Teich}(S)$. In other words, generic directions in $\mathrm{Teich}(S)$ are sublinearly Morse, and almost every sublinearly Morse geodesic is generic.
Our main result shows that sublinear Morseness does not characterize genericity:
Theorem A.
There exist Teichmüller geodesic rays which are sublinearly Morse but have minimal non-uniquely ergodic vertical foliations.
Our examples are on the genus two surface, but they can be lifted to higher genus examples by taking covers. We note that [KM96] proved that tracking rays are recurrent to the thick part of $\mathrm{Teich}(S)$, while Masur’s criterion [Mas82] implies that rays with non-uniquely ergodic vertical foliation are not recurrent. Hence the examples Theorem A exhibit multiple non-generic pathologies.
Our construction utilizes the robust machinery of Chaika–Masur–Wolf [CMW19] for producing Teichmüller geodesic rays with minimal non-uniquely ergodic vertical foliations. We verify sublinear Morseness for our examples via work of Durham–Zalloum [DZ22, Theorem F], which provides a criterion for sublinear Morseness for Teichmüller geodesic rays. To verify the criterion, we utilize hierarchical techniques from [Raf05, Raf14, Dur23] to analyze how quickly the shadows of our geodesic rays diverge in the curve graph $\mathcal{C}(S)$ of $S$. In particular, we show that certain examples from [CMW19] achieve a balance of diverging quickly enough in $\mathcal{C}(S)$ to be sublinearly Morse but not so quickly that they are uniquely ergodic. See [CE07, Tre14, CT17] for the connection between divergence rates and ergodicity.
Theorem A is optimal in the following sense. Sublinear Morseness for a given ray is a property controlled by some sublinear function $\kappa$. The sublinear function in the tracking result from [GQR22] is obtained via a non-constructive argument. However, in [QRT24], Qing–Rafi–Tiozzo show that sample paths of random walks of $\mathcal{MCG}(S)$ on itself track $\log^{p}$-Morse geodesics, for $p=p(S)$ controlled by the topology of the surface $S$. The sublinear function in our example is also a controlled power of $\log$, with the power arising from [DZ22] via a refined version (from [Dur23]) of the hierarchical “passing-up” arguments utilized in [QRT24]. Thus we expect that one cannot skirt around the examples in Theorem A by changing the sublinear function.
We end with an interesting consequence of Theorem A. There is a natural way to associate to any Teichmüller geodesic ray its limit set in $\mathbb{P}\mathcal{MF}(S)$. Notably, Cordes [Cor17] proved that this map is a well-defined injection when restricted to the set of Morse geodesic rays. That is, every Morse geodesic ray determines a unique limit point in $\mathbb{P}\mathcal{MF}(S)$. On the other hand, by Remark 2 and [CMW19, Theorem 2.7], the limit set in $\mathbb{P}\mathcal{MF}(S)$ of the Teichmüller geodesic rays in our examples can be made to be more than a point, including possibly an interval.
Corollary B.
There exist sublinearly-Morse Teichmüller geodesic rays whose limit sets in $\mathbb{P}\mathcal{MF}(S)$ are more than a point.
Hence, even though each sublinearly Morse Teichmüller geodesic determines a unique point in the Gromov boundary of the curve graph by [DZ22, Theorem A], there is no such injection into $\mathbb{P}\mathcal{MF}(S)$ as in [Cor17].
Acknowledgments
We would like to thank Yushan Jiang and Abdul Zalloum for useful discussions. We would particularly like to thank Jon Chaika for providing some useful statements which helped us get started working with [CMW19]. Durham was partially supported by NSF grant DMS-1906487. Wu was partially supported by Simons collaboration grant 850685.
2. Informal discussion of [CMW19]
We begin with an informal description of what is happening in the construction of Chaika–Masur–Wolf [CMW19], which involves in part a detailed analysis of the slit torus examples due to Veech [Vee69].
First, take one copy of the $2$-torus and cut a slit in it. Then lift it to a double cover with two branch points being the endpoints of the slit. This cover is a genus-$2$ surface, $S$. The idea now is to put a particular flat structure on each torus and then deform it. This deformation will lift under the cover to a geodesic in the Teichmüller space for $S$, but it will really live in the subspace corresponding to the Teichmüller space of the twice-punctured torus.
Picking any flat structure on the torus and cutting the slit determines a point in $\mathrm{Teich}(S)$, as well as a quadratic differential. So we want to pick a particular flat structure so that the “vertical” direction is skewed. This is the meaning of the column vectors in the matrix
$\begin{pmatrix}1&-\alpha\\ 0&1\end{pmatrix}.$
This choice of skewed vertical direction determines a lattice in the plane $\mathbb{R}^{2}$. The idea then is to choose a pair of points in the lattice which will correspond to the next vertical and horizontal directions that we want. Flowing along the Teichmüller geodesic flow, we can arrange that the lengths of these new vertical and horizontal lines become comparable to $1$, while the original slit becomes very long.
Now we have a very long slit on the torus, with one endpoint at the corner point (all four corners are identified), while the other point is somewhere in the interior. The last claim follows from arranging the original vertical direction to be skewed at an irrational angle, and also the flow time also being irrational, so that the other endpoint is not at the corner. This is the point of Proposition 1.
With this setup, we can draw an arc from the interior endpoint to the corner, whose length is bounded by the area of the torus. This new slit now lifts to a curve which has bounded length in the corresponding metric on $S$. We repeat this process, flowing to the next pair of vertices, so that the second slit now has very long length with one endpoint at the corner and the other in the interior.
Through the work of [CMW19], we gain explicit control over the slit curves appearing through this process from the information encoded in the continued fraction expansion of $\alpha$. However, we will only need the output of their analysis.
3. Our Teichmüller geodesic and its slit curves
In this section, we show that the construction from [CMW19]. We begin by setting up some notation.
3.1. Setup and notation
In this subsection, we will introduce some of the key players in the paper. We point the reader toward [Wri15] and [FM11] for background on translation surfaces and Teichmüller theory, respectively.
Let $\alpha:=[1,4,9,16,25,...]$ be the real number with given continued fraction expansion. Let $T$ denote the horizontally and vertically oriented square torus and $Y=\begin{pmatrix}1&-\alpha\\ 0&1\end{pmatrix}T$. We glue two isometric, identically oriented copies of $Y$ together along a slit with holonomy $(\sum_{j=1}^{\infty}2(p_{2j}-q_{2j}\alpha),0)$. Denote this flat surface by $X$. Let $g_{t}=\begin{pmatrix}e^{t}&0\\ 0&e^{-t}\end{pmatrix}$, then the action of $g_{t}$ on $X$ induces a Teichmüller geodesic $\gamma$.
Observe that the first return of the vertical flow to the horizontal base is a rotation by $\alpha$. Let $\frac{p_{k}}{q_{k}}=[1,4,9,\dots,k^{2}]$.
Remark 1.
It follows from the definition that $q_{k+1}=(k+1)^{2}q_{k}+q_{k-1}$.
Recall that a measure preserving flow on a space $Y$ with probability measure $m$ is said to be ergodic if $Y$ cannot be written as the disjoint union of two subsets that are invariant under the flow and each of positive measure. The flow is uniquely ergodic if $m$ is the unique invariant measure for the flow.
Lemma 1.
The Teichmüller geodesic $\gamma$ associated with $X$ is non-uniquely ergodic.
Proof.
By [CMW19, Theorem 2.3] (originally done by Veech [Vee69]), if $\alpha=[a_{i}]$ has a subsequence $n_{k}$ such that $\sum_{k=1}^{\infty}a_{n_{k}+1}^{-1}<\infty$, then the associated interval exchange on the surface associated with the ergodic geodesic is non-uniquely ergodic. In our case, since $\alpha=[1,4,9,16,25,...]$, it follows that $\sum_{k=1}^{\infty}a_{k}^{-1}=\sum_{k=1}^{\infty}k^{-2}=\pi^{2}/6<\infty$. ∎
3.2. The geodesic ray and its slit curves
In this subsection, we prove begin our analysis of the geodesic ray $\gamma$. Roughly speaking, we need a sequence of times $\{t_{k}\}$—the gap between which grows logarithmically—so that at each time there is an increasingly short curve (the slit), in the complement of which there are no short curves. In Section 4, we show that this sequence of slit curves diverges in the curve graph $\mathcal{C}(S)$ at a linear rate (in the index of the sequence).
Our first Proposition 1 provides the sequence of times corresponding to the slit curves on the geodesic ray we need, then Proposition 2 shows that the slit curves that are sufficiently far apart are filling, and Proposition 3 in the next subsection providing the combinatorial information we need about how the short curves as those times are organized in $\mathcal{C}(S)$.
Proposition 1.
Let $t_{k}:=\log(q_{2k+1})$,
  1. (1)
    For any fixed $D>0$, there exists a constant $C>0$ so that for every $k$:
    $|t_{k}-t_{k+D}|<C\log t_{k}.$
  2. (2)
    The geodesic flow of $X$ at time $t_{k}$ can be split into two tori $Y^{k}_{\pm}$, each with a pair of slits as the boundary, and both are uniformly $\epsilon$-thick, where $\epsilon$-thick means systolic lengths $\geq\epsilon$.
  3. (3)
    The (hyperbolic) lengths of the splitting slits $\zeta_{k}$ at time $t_{k}$ goes to 00 as $k\to\infty$.
Proof.
Item (1): Since $t_{k}=\log(q_{2k+1})$,
$|t_{k}-t_{k+1}|=|\log(q_{2k+1}/q_{2k+3})|.$
By Remark 1, $q_{k+1}=(k+1)^{2}q_{k}+q_{k-1}$, so
$|t_{k}-t_{k+1}|=\log\left(\frac{(2k+1)^{2}(2k)^{2}q_{2k-1}+q_{2k-1}+(2k+1)^{2}q_{2k-2}}{q_{2k}}\right)=O(\log(k)).$
Moreover, by $|t_{k}-t_{k+D}|\leq|t_{k}-t_{k+1}|+\dots+|t_{k+D-1}-t_{k+D}|$, it follows that
$|t_{k}-t_{k+D}|=O(\log(k)).$
To see that $\log k<\log t_{k}$, recall that $\frac{p_{k}}{q_{k}}$ converges to the continued fraction $\alpha=[1,4,9,16,...]$ (hence each $q_{i}>0$), and recall $t_{k}:=\log(q_{2k+1})$, so $q_{k}=k^{2}q_{k-1}+q_{k-2}>k^{2}q_{k-1}$, so $\frac{q_{k}}{q_{k-1}}>k^{2}$. Therefore, $q_{k}>\prod_{i=1}^{k}i^{2}$, and hence $t_{k}>\log(q_{k})>\sum_{=1}^{k}\log(i^{2})>k$. In particular, $\log(t_{k})>\log(k)$.
Item (2): This follows immediately from Proposition 4.2 of [CMW19].
Item (3): Recall the shear of torus $\alpha$ on the torus, thus on the surface $X$, is given by $\alpha=[a_{i}]$, where $a_{i}=i^{2}$. By [CMW19, Lemma 2.16], if we set $n_{k}=2k+1$,
then the flat length of each slit $\zeta_{k}$ at time $t_{k}$ satisfies
$|\zeta_{k}|\asymp\frac{1}{a_{n_{k}+1}}.$
(Also see the proof of Theorem 2.7 of [CMW19, Section 7.1].) Therefore, $|\zeta_{k}|\to 0,$ as $k\to\infty$. For each slit $\zeta_{k}$, consider a disk $D_{k}$ in the torus containing the $\zeta_{k}$, which is indeed a cylinder. Gluing along the slit of each cylinder gives an annulus. By Riemann Mapping Theorem, this cylinder is conformal to a flat cylinder $C_{k}$.
Now, since flat length $|\zeta_{k}|$ is approaching 00, it implies the the modulus of a cylinder $C_{k}$, say $m_{k}$, would go to infinity. So the extremal length $E(\zeta_{k})$, (by definition, is the reciprocal of the modulus of the biggest cylinder) goes to 00, as $k$ goes to infinity. Now by Maskit’s comparison result [Mas85], we have
$\frac{H(\zeta_{k})}{\pi}\leq E(\zeta_{k}).$
Therefore, the hyperbolic length of the slit $\zeta_{k}$, $H(\zeta_{k})$, goes to 00, as $k$ goes to infinity. ∎
Figure 1. On each torus, find a disk containing the slit; gluing the two disks containing the slit along the slits gives the cylinder. The length of the green segment is the shear of the torus, $\alpha$.
Remark 2.
By Lemma 1, the vertical flow of the flat surface $X$ defined above is non uniquely ergodic. The flat structure of $X$ corresponds to the average of the two ergodic measures ($c=0$ in the language of [CMW19]). If we replace the flat structure with one which correspond to a weighted sum of the two ergodic measure, where weights are non equal but both positive ($c\in(-1,1)$ and $c\not=0$ in the notation of [CMW19]), then for this new flat surface $X_{c}$, the conclusions of Proposition 1 are still valid. In particular:
  • Item (1) only involves the definition of $t_{k}$ and is unrelated to the surface itself.
  • Item (2) can be shown by replacing Proposition 4.2 with Proposition 4.5 of [CMW19] in the proof of Proposition 1.
  • Item (3) is due to the fact that the horizontal component of the slit for $X_{0}=X$ is the average of the horizontal component of the slit for $X_{c}$ and for $X_{-c}$, hence the latter is no more than twice of the slit of $X_{0}$. The vertical components of these slits are all the same. Now apply the same argument as in Proposition 1 to get estimate of the hyperbolic length. Also see the proof of Theorem 2.7 in Section 7 of [CMW19].
To set ourselves up for our combinatorial arguments Section 4, we need the following observation:
Proposition 2.
When $k$ is sufficiently large, $\zeta_{k+3}$ and $\zeta_{k-3}$ fill the surface $X$.
Proof.
As before, denote the geodesic ray by $\gamma$. By [CMW19, Lemma 2.16], when $k$ is large enough, at time $t_{k+3}$, the horizontal length of $\zeta_{k+3}$ $\asymp 1/a_{n_{k+3}+1}=1/a_{2k+8}$, while the vertical length of $\zeta_{k+3}$ $\asymp q_{n_{k+2}}/q_{n_{k+3}}=q_{2k+5}/q_{2k+7}$. So, at time $t_{k+1}$, the vertical length of $\zeta_{k+3}$ would be much larger than $1$, hence would cross both torus vertically many times. In other words, for any point on the geodesic flow of $X$ at time $t_{k+1}$, the shortest horizontal segment from this point to $\zeta_{k+3}$ has length bounded by $O(1)$. We call this length the “horizontal distance”. Hence at time $t_{k}$, the shortest horizontal distance from every point to $\zeta_{k+3}$ $\asymp q_{n_{k}}/q_{n_{k+1}}=O(1/k^{4})$. By similar argument, at time $t_{k}$ the shortest vertical distance from every point to $\zeta_{k-3}$ is bounded by $O(1/k^{4})$. Because the slits are geodesics and the flat surface is non-positively curved, they intersect minimally. If there is some non-trivial loop on the surface which is disjoint from both slits, it has to travel outside a small neighborhood of the slit. However, by the argument above, at time $t_{k}$, outside a small neighborhood of the slit both tori are cut by $\zeta_{k-3}$ and $\zeta_{k+3}$ into a grid of small rectangles of size $O(1/k^{4})$, so such a non-trivial loop is impossible. This shows that $\zeta_{k-3}$ and $\zeta_{k+3}$ fill the surface. ∎
4. Combinatorics of the slit curves
The main goal of this section is Proposition 3 below, in which we prove that the slit curves $\zeta_{k}$ provided by Proposition 1 make linear progress (in $k$) in the curve graph $\mathcal{C}(S)$. See [Min06, MM00] for background on curve graphs and subsurface projections.
Proposition 3.
For any $L>0$, there exists a sequence ${k_{n}}$ and a constant $B=B(S,L)>0$ so that we have
  1. (1)
    $|t_{k_{n}}-t_{k_{n+1}}|<B\log k_{n}$.
  2. (2)
    $d_{\mathcal{C}(S)}(\zeta_{k_{n}},\zeta_{k_{n+1}})>L.$
In Lemma 2 below, we derive a more refined version of Proposition 3, but the proposition above is the key result.
The proof involves an iterated argument which a blend of some results connecting short curves and subsurface projections due to Rafi [Raf05, Raf14] and Modami–Rafi [MR25], and so-called “passing-up” techniques for producing large subsurface projections satisfying certain properties due to Durham [Dur23].
Remark 3.
We note one of the main technical difficulties here is that the slit curves $\zeta_{k}$ need not correspond to boundary curves of subsurfaces with large projections for $\mu_{-},\mu_{+}$. In particular, knowing that they are pairwise filling (Proposition 2) alone is not enough to show that they spread out in $\mathcal{C}(S)$ at a uniform rate.
Proof of Proposition 3.
By Proposition 2, we may assume that the slit curves $\sigma_{i}$ pairwise fill $S$. We fix a subsurface projection threshold constant $L_{0}=L_{0}(S)>0$ to be sufficiently large to make the arguments below work.
For sufficiently large $L_{0}=L_{0}(S)>0$, we can fix $\epsilon=\epsilon(L_{0},S)>0$ to be small enough so that if $\zeta$ is a simple closed curve on $S$ so that $l_{\gamma(t)}(\zeta)<\epsilon$ for some time $t$, then work of Modami–Rafi111[MR25] provides a quantitative version of [Raf05, Theorem 6.1], namely that increasingly short curves produce increasingly large subsurface projections. [MR25] provides a subsurface $V\subset S-\zeta$ (with possibly $V$ the annulus with core $\zeta$) so that $diam_{\mathcal{C}(V)}(\gamma)>L_{0}$.
For the rest of the proof, we will increase $L_{0}$ and $\epsilon$ as necessary while maintaining the dependencies only on $S$.
Consider the sequence of slit curves $\zeta_{k}$. By item (3) of Proposition 1, there exists $K_{0}=K_{0}(\epsilon)>0$ so that if $k>K_{0}$, then $l_{\gamma(t_{k})}(\zeta_{k})<\epsilon$. For each $k>K_{0}$, let $V_{k}$ denote the $L_{0}$-large subsurface in the complement $S-\zeta_{k}$ provided by [Raf05, Theorem 6.1]. We denote the collection of these subsurfaces $V_{k}$ by $\mathcal{V}$.
Our goal is to show that the $\zeta_{k}$ spread out in $\mathcal{C}(S)$ uniformly quickly. Using the $V_{k}$ as surrogates, the following claim forces the boundary curves of any sufficiently large subcollection to spread out in some subsurface, with extra control over where the slit curves lie:
Claim 1.
For any $L_{1}>L_{0}$, there exists $N_{1}=N_{1}(S,L_{0},L_{1})>0$ so that for any collection $\mathcal{U}\subset\mathcal{V}$ with #$\#\mathcal{U}\geq N_{1}$, there exists a subsurface $W$ and $V_{j},V_{j},V_{k},V_{l}\subset W$ with $V_{i},V_{j},V_{k},V_{l}\in\mathcal{U}$ so that
  1. (1)
    $d_{\mathcal{C}(W)}(\mu^{-},\mu^{+})>L_{1}$, and
  2. (2)
    All pairwise distances in $\mathcal{C}(W)$ between $\mu_{-},\mu_{+},\partial V_{i},\partial V_{j},\partial V_{k},\partial V_{l}$ are larger than $\frac{L_{1}}{100}$, and $\partial V_{i},\partial V_{j},\partial V_{k},\partial V_{l}$ appear in that order along any geodesic in $\mathcal{C}(W)$ between $\mu_{-},\mu_{+}$ (see Figure 2).
  3. (3)
    At least one of the slit curves $\zeta_{i},\zeta_{j},\zeta_{k},\zeta_{l}$ corresponding to $V_{i},V_{j},V_{k},V_{l}$ is contained in $W$.
[…]
Figure 2. Using [Dur23, Proposition 4.7], we can arrange for the boundary curves $\partial V_{i},\partial V_{j},\partial V_{k},\partial V_{l}$, and hence $\zeta_{i},\zeta_{j},\zeta_{k},\zeta_{l}$, appear in order along any geodesic in $\mathcal{C}(W)$ between $\mu_{-},\mu_{+}$.
Proof of Claim 1.
Items (1) and (2) are essentially an immediate consequence of [Dur23, Proposition 4.7], which not only provides a large domain $W$ containing some of the $V_{k},\dots,V_{k+N_{1}}$ used to produce it, but forces some of their boundary curves to roughly evenly distribute along the geodesic in $\mathcal{C}(W)$ between $\pi_{W}(\mu_{-})$ and $\pi_{W}(\mu_{+})$. In particular, by choosing the subdivision constant to be $\sigma=\frac{1}{100}$ as in the statement of [Dur23, Proposition 4.7], we can produce $V_{i},V_{j}$ satisfying the desired conclusions.
For item (3), first observe that since the $\zeta_{i},\zeta_{j},\zeta_{k},\zeta_{l}$ pairwise fill $S$ and each of $V_{i},V_{j},V_{k},V_{l}\subset W\subset S$, it is not possible for any of the $\zeta_{i},\zeta_{j},\zeta_{k},\zeta_{l}$ to be disjoint from $W$, for if $\zeta_{i}$ were disjoint from $W$, then $\zeta_{j}$ would have to intersect $\partial V_{j}$, which is impossible. Hence they either must intersect $\partial W$, be contained in $\partial W$, or be contained in $W$ itself. Moreover, at most one of the slit curves can be contained in $\partial W$, so we may assume, up to reindexing, that $\zeta_{i},\zeta_{j},\zeta_{k}$ are not contained in $\partial W$.
By [Raf14, Theorem 5.3], there is an active interval of times $I_{V_{i}}$ along the Teichmüller geodesic $\gamma$ during which the curves in $\partial V_{i}$ are shorter than $\epsilon$ and outside of which $\gamma$ has a bounded diameter projection to $\mathcal{C}(V_{i})$, and similarly for $V_{j},V_{k},W$. By our choice of the arrangement and location of $\partial V_{i},\partial V_{j},\partial V_{k}$ along any geodesic in $\mathcal{C}(W)$ between $\mu_{-},\mu_{+}$ and the fact that the projection of $\gamma$ to $\mathcal{C}(W)$ is a uniform (unparameterized) quasigeodesic [Raf14, Theorem 6.1], we may increase $L_{0}=L_{0}(S)>0$ as necessary to arrange that $I_{V_{i}},I_{V_{j}},I_{V_{k}}\subset I_{W}$ and that these intervals appear in this order along the parameter interval $[0,\infty)$ for $\gamma$.
Finally, observe that if $t\in[0,\infty)$ is such that $l_{\gamma(t)}(\zeta_{j})<\epsilon$, then $t$ comes after $I_{V_{i}}$ and before $I_{V_{k}}$, again because of [Raf14, Theorem 6.1]. Hence $t\in I_{W}$, and in particular $\zeta_{j}$ and $\partial W$ are short simultaneously along $\gamma$. Since this is only possible when $\zeta_{j}$ is contained in $W$ by the Collar Lemma, this proves item (3) of the claim, as required. ∎
The rest of the argument proceeds by induction on the the complexity of the subsurfaces produced using Claim 1 and its analogues below. For any surface $Y$, recall that its topological complexity is $\xi(Y)=g(Y)+p(Y)$, where $g(Y)$ counts its genus and $p(Y)$ counts its punctures. Note that when $X\subset Y$ is a proper subsurface, we have $\xi(X)<\xi(Y)$.
We also require some more organizational notation. Since the domains in $\mathcal{V}$ come in an order provided by the order of the slit curves $\zeta_{i}$, we can partition $\mathcal{V}$ into subcollections $\mathcal{V}_{j}$ of consecutive domains, with each of the #$\#V_{j}=N_{1}$ (for $N_{1}$ as in Claim 1), and the domains in $\mathcal{V}_{j}$ immediately preceding those in $\mathcal{V}_{j+1}$ for every $j\geq 1$.
For each $j$, Claim 1 provides potentially many subsurfaces $W$ satisfying the conclusions of the claim for $\mathcal{U}=\mathcal{V}_{j}$. Choose $W_{j}$ to be one of the subsurfaces that maximizes topological complexity among all such subsurfaces, and let $\mathcal{W}$ be the collection of these $W_{j}$. Given $W\in\mathcal{W}$, let $J_{W}=\{j|W_{j}=W\}$.
Claim 2.
For any $W\in\mathcal{W}$ with $W\neq S$, we have #$\#J_{W}=1$.
Proof of Claim 2.
Suppose that $W=W_{i}$ and $W=W_{j}$ for $i\neq j$. Then item (4) of Claim 1 provides distinct slit curves $\zeta_{i}\neq\zeta_{j}$ with $\zeta_{i},\zeta_{j}\subset W$, which is impossible since $\zeta_{i},\zeta_{j}$ fill all of $S$. This completes the proof. ∎
Claim 2 says that the $W_{i}$ are all distinct and, moreover, they have topological complexity at least $3$.
Claim 3.
The conclusions of Claim 1 hold when replacing the collection $\mathcal{V}$ with the collection $\mathcal{W}$.
Proof.
Since items (1) and (2) follow from the generalities in [Dur23, Proposition 4.7], the only item to check is item (3). However, for each $i$, we have the containment $\zeta_{i},V_{i}\subset W_{i}$, so any container domain $S$ produced by [Dur23, Proposition 4.7] necessarily satisfies $\zeta_{j},V_{j},\subset W_{j}\subset X$ for any of the four indices produced in the claim. This completes the proof. ∎
Once again partitioning the domains in $\mathcal{W}$ as we did with $\mathcal{V}$ above into ordered subcollections, we can use Claim 3 to produce another group $\mathcal{Z}$ of container domains for the $W_{i}$ which are of complexity at least $4$. Moreover, it only takes $N_{1}$-many of the $W_{i}\in\mathcal{W}$ to produce each such container, while each such $W_{i}$ only took $N_{1}$-many of the $V_{j}\in\mathcal{V}$ to produce.
Since $\xi(S)=6$, we can repeat this process at most two more times, at each level only increases the number of consecutive domains in $\mathcal{V}$ that we need by a multiplicative factor of $N_{1}=N_{1}(S)$. In particular, by taking any collection $\mathcal{A}$ of $N_{1}^{6}$-many consecutive domains from $\mathcal{V}$, we can use the appropriate version of Claim 1 to produce domains $Z_{i},Z_{j}$ so that
  • There are slit curves $\zeta_{i},\zeta_{j}$ and $V_{i},V_{j}$ so that $V_{i}\subseteq Z_{i}$ and $V_{j}\subseteq Z_{j}$, with either $\zeta_{i}\subset V_{i}$ (if $V_{i}\subsetneq Z_{i}$) or $\zeta_{i}\perp V_{i}$ (if $V_{i}=Z_{i}$), and similarly for $j$;
  • $d_{S}(\partial Z_{i},\partial Z_{j})>\frac{L_{0}}{100}$.
Then since the slit curves $\zeta_{i},\zeta_{j}$ project at most distance $1$ from $\partial Z_{i},\partial Z_{j}$ in $\mathcal{C}(S)$, we see that $d_{S}(\zeta_{i},\zeta_{j})>\frac{L_{0}}{100}-2$. On the other hand, since the slit curves become short along $\gamma$ in their given order and $\gamma$ projects to a uniform (unparameterized) quasigeodesic in $\mathcal{C}(S)$, we have that the slit curves corresponding to first and last domains in $\mathcal{A}$ must also project roughly $\frac{L_{0}}{100}$ apart in $\mathcal{C}(S)$. This completes the proof of item (2) of Proposition 3, and item (1) of this proposition follows from item (1) of Proposition 1.
In the following lemma, we record the information that we actually need from Proposition 3. Roughly, the lemma says that we can find a sequence of times $\{t_{k}\}$, whose spacing as a time parameter grows like $\log$ (in the index $k$), but whose spacing in the curve graph grows linearly (with the index $k$), with the corresponding intervals in the curve graph having bounded overlap.
We set some notation for the statement. Recall that Masur–Minsky [MM99] proved that Teichmüller geodesics project to unparameterized quasi-geodesics with uniform constants controlled by (the topology of) $S$. Since $\gamma$ has a minimal non-uniquely ergodic vertical foliation, its projection $\pi_{S}(\gamma)\subset\mathcal{C}(S)$ is a quasi-geodesic ray. Fix a geodesic ray $\Gamma$ in its asymptotic class [^2] Despite the fact that $\mathcal{C}(S)$ is locally infinite, one can find geodesic ray representatives using the finiteness properties of tight geodesics [Bow08].. For each $k$, let
$\Gamma_{k}=\pi_{\Gamma}(\pi_{S}(\gamma|_{[t_{k},t_{k+1}]}))$
denote the closest-point projection to $\Gamma$ in $\mathcal{C}(S)$ of the projection to $\mathcal{C}(S)$ of the restriction of $\gamma$ to $[t_{k},t_{k+1}]$.
Lemma 2.
For our Teichmüller geodesic $\gamma$ and any sufficiently large $L=L(S)>0$, there exists a sequence of times $\{t_{k}\}$ and constants $C=C(S)>0$ so that for all $k$, we have
  1. (1)
    $|t_{k}-t_{k+1}|<C\log k$;
  2. (2)
    $d_{\mathcal{C}(S)}(\gamma(t_{k}),\gamma(t_{k+1}))>L$;
  3. (3)
    $\mathrm{diam}_{\mathcal{C}(S)}(\Gamma_{k})>L/2$;
  4. (4)
    If $j=k-1$ or $j=k+1$ then $\mathrm{diam}_{\mathcal{C}(S)}(\Gamma_{k}\cap\Gamma_{j})<L/10$;
  5. (5)
    If $j\neq k-1,k,k+1$, then $\Gamma_{j}\cap\Gamma_{k}=\emptyset$.
Proof.
Fixing $L=L(S)>0$ below, item (1) follows easily from the proof of part 1 of Proposition 1. Item (2) follows from Proposition 3. Item (3) follows from item (2) plus the fact that $\pi_{S}(\gamma)$ is a uniform (in $S$) unparameterized quasi-geodesic, and hence is uniformly close (in $S$) to $\Gamma$ by uniform hyperbolicity of $\mathcal{C}(S)$.
For item (4), since $\pi_{S}(\gamma)$ is a uniform (unparameterized) quasi-geodesic, it can backtrack at most a bounded amount. By hyperbolicity of $\mathcal{C}(S)$, this backtracking is efficiently recorded on the geodesic ray $\Gamma$. In particular, by choosing $L=L(S)>0$ sufficiently large, we can guarantee that the overlap of $\Gamma_{k}$ with $\Gamma_{k-1}$ or $\Gamma_{k+1}$ is a controlled fraction of $L$.
Item (5) now follows easily from items (3) and (4), completing the proof. ∎
5. Confirming sublinear Morseness
In this final section, we confirm that our examples are sublinearly Morse. For our purposes, the definition of sublinearly Morse is not strictly necessary; see [QRT24, DZ22]. Instead, we use the following criterion from [DZ22] for a Teichmüller geodesic to be sublinearly-Morse.
Definition 1.
For a given sublinear function $\kappa$, we say that a Teichmüller geodesic $\beta$ satisfies the $\kappa$-bounded projections property if
$d_{Y}(\beta(0),\beta(t))\leq C\cdot\kappa(t),$
for some constant $C=C(S)>0$ and for all proper subsurfaces $Y\subsetneq S$.
The following is [DZ22, Theorem K, part 2]:
Theorem 1.
There exists a $p=p(S)>0$ so that if $\kappa^{2p}$ is sublinear and $\beta$ is a Teichmüller geodesic with $\kappa$-bounded projections, then $\beta$ is $\kappa^{2p}$-Morse.
The following completes the proof of Theorem A:
Theorem 2.
There exists $p=p(S)$ so that the Teichmüller geodesic $\gamma$ associated with $X$ is $\log^{2p}$-Morse.
Proof.
By Theorem 1, it suffices to prove that $\gamma$ has $\log$-bounded projections. Fix a subsurface $Y\subset S$.
As in Lemma 2 above, let $\Gamma$ be a geodesic representative in $\mathcal{C}(S)$ of the quasi-geodesic ray $\pi_{S}(\gamma)$. By the Bounded Geodesic Image Theorem [MM00, Theorem 3.1], if $\partial Y$ is not close to $\Gamma$, then $\gamma$ has bounded projections to $\mathcal{C}(Y)$.
On the other hand, the closest point projection $P_{Y}$ of $\partial Y$ to $\Gamma$ has uniformly bounded diameter. By increasing the constant $L$ in Lemma 2 a bounded amount as necessary, we can guarantee that $P_{Y}$ overlaps at most two consecutive $\Gamma_{k}$, and hence that $\partial Y$ is as far as we would like from the restrictions $\gamma_{k}=\pi_{S}(\gamma|_{[t_{k},t_{k+1}]})$ for all but two consecutive $k$.
In particular, this forces the active interval $I_{Y}$ to overlap at most two consecutive $[t_{k},t_{k+1}]$, while on the other hand the projection to $\mathcal{C}(Y)$ is coarsely constant outside of the active interval $I_{Y}$ for $Y$ by [Raf14, Theorem 6.1].
Supposing that $I_{Y}$ is contained in the union of $[t_{k-1},t_{k}]\cup[t_{k},t_{k+1}]$, then the distance formula for $\mathrm{Teich}(S)$ [Raf07, Dur16] and [Raf14, Theorem 6.1] provide that:
(1) $d_{Y}(\gamma(0),\gamma(t))$ $\asymp$ $d_{Y}(\gamma(t_{k-1}),\gamma(t))$
(2) $\prec$ $d_{Y}(\gamma(t_{k-1}),\gamma(t_{k+1}))$
(3) $\prec$ $\sum_{W}d_{W}(\gamma(t_{k-1}),\gamma(t_{k+1}))$
(4) $\asymp$ $d_{\mathrm{Teich}(S)}(\gamma(t_{k-1}),\gamma(t_{k+1}))$
(5) $\prec$ $2\log k$
(6) $\prec$ $2\log t_{k}$
(7) $\leq$ $2\log t.$
In these computations, the coarse (in)equalities $\asymp$ and $\prec$ are (in)equalities which hold up to bounded additive and multiplicative constants depending only on the topology of $S$. Moreover, when $Y$ is an annulus, then the distance appearing in both lines (1) and (2) and the corresponding term in line (3) are all distances in the combinatorial horoball over $\mathcal{C}(Y)$ (which takes the length of the core curve of $Y$ into account), but this does not affect the computation. Finally, the penultimate inequality follows from item (1) of Lemma 2, where we replace $t_{k}$ with $t_{k-1}$ in line (5) depending (respectively) on whether $t\geq t_{k}$ or $t<t_{k}$.
In particular, $\gamma$ as $\log$-bounded projections, as required. ∎